OpenStudy (anonymous):

United Parcel Service has contracted you to design an open box with a square base that has a volume of 5000 cubic inches. A.) Express the surface area S of the box as a function of x?

OpenStudy (campbell_st):

let the length of 1 side of the base be x height = h then the volume V = x^2h using the given information then 5000 = x^2h let h = 5000/x^2 Surface area... not top  the box has dimensions x, x, h base = x^2 4 sides all have same area xh so surface area s = x^2 + 4xh use the substitution h = 5000/x^2 then s = x^2 + 4x(5000/x^2) or s = x^2 + 20000/x